# A Systematic Search for Spectral Hardening in Blazar Flares with the Fermi-Large Area Telescope

### v1

@@ 3 Methodology
We investigate the presence of spectral hardening during flaring episodes. These flaring episodes, characterized by increased activity, are selected from the LCs and are expected to enhance the probability of detecting the spectral hardening feature. The reason is that synchrotron peaks may shift towards higher energies during flares, and blazar spectra become harder in the -ray band when these sources are flaring (e.g., Archambault et al. 2015; Ahnen et al. 2018; Valverde et al. 2020; Acciari et al. 2020; Sahu et al. 2021; Abe et al. 2024). Secondly, our data analysis pipeline is run over the flaring periods, leading to a TShardening profile, which we use for evaluating the significance and robustness of the spectral hardening feature.
@@ 3.1 Flare Selection
To select flares in the LCs, we employ a systematic approach developed by Wagner et al. (2022). This method characterizes and models the variability of the LC using Bayesian Block Analysis333https://docs.astropy.org/en/stable/api/astropy.stats.bayesian_blocks.html (BBA) as described by Scargle et al. (2013). BBA is a non-physical model designed to detect statistically significant variations in the data while minimizing the influence of observational or random uncertainties. Figure 1 provides a visual representation of this procedure.
@@
BBA divides the data into blocks by considering their variations, with the granularity of these blocks controlled by parameters such as (Scargle et al. 2013). For our analysis, we adopt , which reflects a standard balance between detecting true variability in blazar LCs and minimizing false detections444https://github.com/swagner-astro/lightcurves?tab=readme-ov-file. The number of detected flares depends on both the choice of and the number of time bins in the LC, as these parameters influence the segmentation of the BBA algorithm. The corresponding value, which is another relevant parameter in BBA, is not fixed but is dynamically determined within the code, following the formulation by Scargle et al. (2013). This ensures that the block segmentation remains adaptive to the characteristics of each individual LC, and while variations in may lead to different flare counts, the adopted selection criteria ensure the robustness of our results. We only consider for the BBA, fluxes in time bins of the LC with TS ¿ 1, which roughly corresponds to a detection significance555For reference, TS values of 4, 9, and 25 correspond approximately to 2, 3, and 5 significance levels, respectively, assuming a distribution with one degree of freedom.. While this threshold ensures a basic level of statistical significance in the LCs, we acknowledge that this criterion does not necessarily eliminate all outliers (Abdollahi et al. 2023). A more refined approach, such as filtering based on the vs. TS relation, could improve the robustness of the selection process and represents a possible refinement for future analyses. However, since the most significant flaring periods tend to have high TS values, the difference between both approaches is not expected to substantially alter the overall results. All settings follow those described by Wagner et al. (2022).
@@
We use two approaches to define a flare: the Baseline and Flip methods. Both are based on the HOP algorithm (Eisenstein & Hut 1998), which selects peaks as the highest local maxima of a flare, with the flare itself referred to as a HOP group. Bayesian blocks between two flares are defined as valley blocks. Throughout this work, we use the term flare to denote periods of increased -ray activity, which in our analysis typically span from several months to over a year, reflecting the near monthly binning of the LCs and the long-term variability patterns in blazars. Finally, to ensure the reliability of the flare identification, we require that each selected flare contains fewer than 50% UL bins. This condition guarantees that the majority of the time bins are actual detections, which is essential for characterizing the flare structure and enabling robust spectral analysis. As a result of this filtering, the final sample includes only well-defined flares with reliable temporal profiles, suitable for a systematic spectral study.
@@ 3.1.1 Baseline Method
The Baseline method, introduced by Meyer et al. (2019), defines a constant flux level as the baseline to select the onset of a flare. If the flux exceeds this baseline, the time period is categorized as a flare. Only blocks with flux values above the baseline are recognized as flares, and adjacent blocks surrounding a peak block that remain above the baseline are considered part of the same flare event within the limitations of the binning chosen. We use the mean flux of each LC as the baseline, consistent with the approach in Wagner et al. (2022). While this choice may result in fewer selected flares compared to using the quiescent or low-state flux, it provides a conservative approach that minimizes the inclusion of spurious events. Figure 1 visually demonstrates the implementation of the Baseline method. While this method relies on the assumption of a baseline, it is a straightforward and effective approach for flare selection. This method identifies a total of 2193 flares; after applying the 50% UL within-the-flare criterion, 1660 flares remain.
@@ 3.2 Spectral Hardening Analysis Framework
After running our data analysis pipeline over all flaring periods selected by the Baseline and Flip methods, this is, a total of 4063 flares, we consider as candidate events those with TS. Since the BPL model introduces two additional degrees of freedom, TShardening is expected to follow a distribution with two degrees of freedom under the null hypothesis. We have verified this assumption by comparing the empirical distribution of TShardening with the theoretical expectation and found good agreement. Given our sample of 4063 flares and the chosen threshold of TS, we expect approximately 10 false positives under the null hypothesis, based on the cumulative probability of the distribution with two degrees of freedom. For higher thresholds, such as TS, the expected number of false positives across the entire sample drops below one, providing a useful benchmark for identifying exceptionally rare events. In this context, a TShardening value of 12 corresponds approximately to a 3 confidence level, and TS to beyond 3.5, offering a familiar interpretation of the strength of the spectral deviations.
The three events with TS undergo meticulous individual examination to complement the automatic safety checks that are applied to all flares, in order to: (1) manually verify the convergence of the fits, since rare cases of short integration times or limited statistics can still cause convergence issues even with the pipeline’s internal checks; (2) inspect the TShardening profile as a function of break energy, which should be smooth with a clear maximum; and (3) check for nearby LAT sources that may contaminate the spectrum. We analyze whether -ray sources within the ROI could affect the result by examining whether any sources other than the target show varying parameters during spectral fitting. As an additional test, not part of the main analysis pipeline, we rerun the pipeline with these nearby sources fixed to assess whether the detection of spectral hardening is robust to potential contamination. This check is used to confirm that the result is not driven by variability in neighboring sources, but the default analysis keeps such sources free to ensure accurate error estimation. All these targeted checks return positive results, confirming that the identified spectral hardening features are robust and not artifacts of the fitting process or source confusion.
We test the stability of the spectral indices for target blazars that have a bright nearby source (TS ¿ 100) flaring simultaneously, focusing on cases where the target source’s flare lasts longer than that of the nearby source. In these cases, the spectral indices remain consistent when comparing the time range in which only the target source is flaring to the full duration of the flare.
@@ 4.1 Spectral Hardening in Blazar Flares
We identify three flares with indications of spectral hardening, all of which fall near or just above our detection threshold (see Table 1). One flare is associated with 4FGL J0238.43116 (1RXS J023832.6311658) with TS, and two with 4FGL J2158.83013 (PKS 2155304), detected independently by the Flip (TS) and Baseline (TS) methods. The flares in PKS 2155304 overlap in time and likely correspond to the same physical event, captured slightly differently due to the distinct selection criteria of the two methods.
@@
Both blazars have been detected by Imaging Atmospheric Cherenkov Telescopes (IACTs): J0238.43116 (Gaté et al. 2017) and PKS 2155304 (Abramowski et al. 2010). The latter has been extensively studied in searches for quasi-periodic oscillations (QPOs; Peñil et al. 2020; Rueda et al. 2022; Peñil et al. 2022; Rico et al. 2025), making it a particularly compelling target for investigating a possible link between spectral hardening and QPOs. Such a connection could help characterize periodic modulations in the jet’s particle acceleration and cooling processes.
@@ 4.2 Physical Interpretation
The observed spectral hardening in the -ray spectrum may indicate a shift of the synchrotron and IC peaks to higher frequencies in the context of a one-zone leptonic SSC model. This interpretation has been proposed in previous works based on individual flaring episodes, such as in the source 1ES 0502+675, where a spectral hardening feature was observed around 1 GeV (Zeng et al. 2022), although we do not reproduce this result. In their one-zone SSC model, the authors report a low magnetic field ( Gauss) and a high minimum Lorentz factor (). Such conditions are expected for TeV blazars, where a low magnetic field reduces synchrotron cooling, allowing emissions to reach very high energies (e.g., Tavecchio et al. 2009; Lefa et al. 2011; Costamante et al. 2018; Biteau et al. 2020). However, the one-zone SSC model, while often used for its simplicity, may struggle to reproduce all observational features and might require extreme physical parameters far from equipartition (e.g., Sikora et al. 1994; Dmytriiev et al. 2021; Nievas Rosillo et al. 2022).
Another proposed scenario involves a second SSC component becoming dominant at higher energies (Abdo et al. 2010a). This would imply that the spectral hardening arises from the transition between distinct emission regions within the jet, each characterized by different physical conditions such as magnetic field strength or electron distribution. The limitations of one-zone models for extreme blazar states support the idea of multiple emission regions. This has been illustrated in previous flares, such as those of PKS 2155304 in 2006 (Abramowski et al. 2012) and Mrk 421 in 2010 (Dmytriiev et al. 2021), where successful modeling required a combination of a steady and a compact emission region. A more recent two-zone model applied to Mrk 501 also reproduces shifting synchrotron peaks during different epochs (Abe et al. 2024).
Alternatively, the hardening may be related to a freshly accelerated electron distribution produced by magnetic reconnection, yielding a very hard synchrotron component. In this case, the feature could arise from a two-population electron distribution within the same emission zone, or from a distinct synchrotron component in a separate region. Similar interpretations have been discussed for the EHBL-like state of Mrk 501 (Sahu et al. 2025), which exhibits rapid spectral evolution and variability difficult to reconcile with one-zone models.
In hadronic scenarios, a proton synchrotron origin would require stronger magnetic fields, lower minimum Lorentz factors, and smaller Doppler factors to produce TeV emission. These models can lead to jet powers approaching the Eddington limit (e.g., Biteau et al. 2020), often in tension with the low disk luminosities observed (e.g., Dmytriiev et al. 2021). However, mixed leptonic-hadronic models involving hadronic cascades in the MeV–GeV range have been proposed, especially for blazars that may be neutrino emitters. A two-zone photohadronic model has also been invoked to explain TeV flaring in Mrk 501 during 2014 (Sahu et al. 2025), supporting a possible hadronic role in VHE emission.
For FSRQs, MeV hardening features might result from thermal emission from the disk or broad-line region (Wierzcholska & Wagner 2025). However, this is unlikely to apply to the EHSP BL Lacs considered here, which typically lack strong external photon fields and have radiatively inefficient disks. In these cases, intrinsic jet processes remain the most plausible explanation.
@@
Our pipeline analyzed 4063 flares and identified two independent flaring episodes with indications of spectral hardening: one in 4FGL J0238.43116 and another in PKS 2155304. The latter is detected with both the Flip and Baseline methods but corresponds to the same flaring activity period. Since both events lie at or just above the detection threshold (TS), the occurrence rate of candidate spectral hardening events is below 0.1%. To our knowledge, there are no theoretical works that provide explicit predictions for the occurrence rate of spectral hardening flares in blazars, as most studies focus on modeling individual events under assumed physical conditions rather than estimating their population frequency (e.g., Tavecchio et al. 2009; Böttcher et al. 2013). This confirms that such features are extremely rare in -ray blazars, if present at all. In the context of standard leptonic models, this rarity suggests that the physical conditions required to produce sharp spectral breaks, such as transitions between distinct electron populations or emission zones, are seldom met. Alternatively, the hardening may emerge only under exceptional jet environments involving magnetic reconnection, multi-zone structures, or hadronic contributions. The scarcity of events reinforces the notion that the dominant blazar emission mechanism is well described by smoothly varying power-law spectra across the Fermi-LAT range, with sharp hardenings representing rare deviations from this behavior.
@@ 5 Summary and Conclusions
One of these events is found in 4FGL J0238.43116 and the other two in PKS 2155304, the latter likely corresponding to the same physical flare captured by both selection methods. Given the small number of detections and their moderate significance, these findings are statistically compatible with random fluctuations and do not constitute strong evidence for a widespread occurrence of spectral hardening among blazars. Nonetheless, the sources identified remain interesting for follow-up studies, particularly due to their known high synchrotron peak frequencies and previous detections at very high energies. Notably, both J0238.43116 and PKS 2155304 have been detected by IACTs. The latter, as a well-studied QPO candidate, stands out as a key target for investigating a possible connection between spectral hardening and QPOs.
Future multi-wavelength observations are essential to further explore these sources, particularly involving IACTs for VHE -ray studies. Our work lays the groundwork for these efforts by identifying key targets and characterizing their tentative spectral hardening behavior. Such observations will be crucial in refining our understanding of the physical processes at play in these extreme blazars and in clarifying the origins of the observed spectral features.
@@ Acknowledgements.
The authors are thankful to D. Yan for private communication on the 1ES 0502+675 analysis. We thank J. Valverde for her helpful revision of the manuscript and J. Ballet, Filippo D’Ammando, and Miguel Ángel Sánchez-Conde for fruitful comments. A. Dinesh and J.L.C. acknowledge the support of MCIN project PID2022-138132NB-C42. A. Domínguez is thankful for the support of the project PID2021-126536OA-I00 funded by MCIN/AEI/10.13039/501100011033.The Fermi-LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis. These include the National Aeronautics and Space Administration and the Department of Energy in the United States, the Commissariat à l’Energie Atomique and the Centre National de la Recherche Scientifique / Institut National de Physique Nucléaire et de Physique des Particules in France, the Agenzia Spaziale Italiana and the Istituto Nazionale di Fisica Nucleare in Italy, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), High Energy Accelerator Research Organization (KEK) and Japan Aerospace Exploration Agency (JAXA) in Japan, and the K. A. Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden. Additional support for science analysis during the operations phase is gratefully acknowledged from the Istituto Nazionale di Astrofisica in Italy and the Centre National d’Études Spatiales in France. This work was performed in part under DOE Contract DE-AC02-76SF00515.

### v2

@@ 3 Methodology
We investigated the presence of spectral hardening during flaring episodes. These flaring episodes, characterized by increased activity, are selected from the LCs and are expected to enhance the probability of detecting the spectral hardening feature. The reason is that synchrotron peaks may shift toward higher energies during flares, and blazar spectra become harder in the -ray band when these sources are flaring (e.g., Archambault et al. 2015; Ahnen et al. 2018; Valverde et al. 2020; Acciari et al. 2020; Sahu et al. 2021; Abe et al. 2024). Secondly, our data analysis pipeline is run over the flaring periods, leading to a TShardening profile, which we used for evaluating the significance and robustness of the spectral hardening feature.
@@ 3.1 Flare selection
To select flares in the LCs, we employed a systematic approach developed by Wagner et al. (2022). This method characterizes and models the variability of the LC using Bayesian Block Analysis333https://docs.astropy.org/en/stable/api/astropy.stats.bayesian_blocks.html (BBA) as described by Scargle et al. (2013). BBA is a non-physical model designed to detect statistically significant variations in the data while minimizing the influence of observational or random uncertainties. Figure 1 provides a visual representation of this procedure.
@@
BBA divides the data into blocks by considering their variations, with the granularity of these blocks controlled by parameters such as (Scargle et al. 2013). For our analysis, we adopted , which reflects a standard balance between detecting true variability in blazar LCs and minimizing false detections444https://github.com/swagner-astro/lightcurves?tab=readme-ov-file. The number of detected flares depends on both the choice of and the number of time bins in the LC, as these parameters influence the segmentation of the BBA algorithm. The corresponding value, which is another relevant parameter in BBA, is not fixed but is dynamically determined within the code, following the formulation by Scargle et al. (2013). This ensures that the block segmentation remains adaptive to the characteristics of each individual LC, and while variations in may lead to different flare counts, the adopted selection criteria ensure the robustness of our results. We only considered for the BBA, fluxes in time bins of the LC with TS ¿ 1, which roughly corresponds to a detection significance555For reference, TS values of 4, 9, and 25 correspond approximately to 2, 3, and 5 significance levels, respectively, assuming a distribution with one degree of freedom.. While this threshold ensures a basic level of statistical significance in the LCs, we acknowledge that this criterion does not necessarily eliminate all outliers (Abdollahi et al. 2023). A more refined approach, such as filtering based on the vs. TS relation, could improve the robustness of the selection process and represents a possible refinement for future analyses. However, since the most significant flaring periods tend to have high TS values, the difference between both approaches is not expected to substantially alter the overall results. All settings follow those described by Wagner et al. (2022).
@@
We used two approaches to define a flare: the Baseline and Flip methods. Both are based on the HOP algorithm (Eisenstein & Hut 1998), which selects peaks as the highest local maxima of a flare, with the flare itself referred to as a HOP group. Bayesian blocks between two flares are defined as valley blocks. Throughout this work, we use the term flare to denote periods of increased -ray activity, which in our analysis typically span from several months to over a year, reflecting the near-monthly binning of the LCs and the long-term variability patterns in blazars. Finally, to ensure the reliability of the flare identification, we require that each selected flare contains fewer than 50% UL bins. This condition guarantees that the majority of the time bins are actual detections, which is essential for characterizing the flare structure and enabling robust spectral analysis. As a result of this filtering, the final sample includes only well-defined flares with reliable temporal profiles, suitable for a systematic spectral study.
@@ 3.1.1 Baseline method
The Baseline method, introduced by Meyer et al. (2019), defines a constant flux level as the baseline to select the onset of a flare. If the flux exceeds this baseline, the time period is categorized as a flare. Only blocks with flux values above the baseline are recognized as flares, and adjacent blocks surrounding a peak block that remain above the baseline are considered part of the same flare event within the limitations of the binning chosen. We used the mean flux of each LC as the baseline, consistent with the approach in Wagner et al. (2022). While this choice may result in fewer selected flares compared to using the quiescent or low-state flux, it provides a conservative approach that minimizes the inclusion of spurious events. Figure 1 visually demonstrates the implementation of the Baseline method. While this method relies on the assumption of a baseline, it is a straightforward and effective approach for flare selection. This method identifies a total of 2193 flares; after applying the 50% UL within-the-flare criterion, 1660 flares remain.
@@ 3.2 Spectral hardening analysis framework
After running our data analysis pipeline over all flaring periods selected by the Baseline and Flip methods, this is, a total of 4063 flares, we considered as candidate events those with TS. Since the BPL model introduces two additional degrees of freedom, TShardening is expected to follow a distribution with two degrees of freedom under the null hypothesis. We verified this assumption by comparing the empirical distribution of TShardening with the theoretical expectation and found good agreement. Given our sample of 4063 flares and the chosen threshold of TS, we expect approximately 10 false positives under the null hypothesis, based on the cumulative probability of the distribution with two degrees of freedom. For higher thresholds, such as TS, the expected number of false positives across the entire sample drops below one, providing a useful benchmark for identifying exceptionally rare events. In this context, a TShardening value of 12 corresponds approximately to a 3 confidence level, and TS to beyond 3.5, offering a familiar interpretation of the strength of the spectral deviations.
The three events with TS undergo meticulous individual examination to complement the automatic safety checks that are applied to all flares, in order to: (1) manually verify the convergence of the fits, since rare cases of short integration times or limited statistics can still cause convergence issues even with the pipeline’s internal checks; (2) inspect the TShardening profile as a function of break energy, which should be smooth with a clear maximum; and (3) check for nearby LAT sources that may contaminate the spectrum. We analyzed whether -ray sources within the ROI could affect the result by examining whether any sources other than the target show varying parameters during spectral fitting. As an additional test, not part of the main analysis pipeline, we rerun the pipeline with these nearby sources fixed to assess whether the detection of spectral hardening is robust to potential contamination. This check is used to confirm that the result is not driven by variability in neighboring sources, but the default analysis keeps such sources free to ensure accurate error estimation. All these targeted checks return positive results, confirming that the identified spectral hardening features are robust and not artifacts of the fitting process or source confusion.
We tested the stability of the spectral indices for target blazars that have a bright nearby source (TS ¿ 100) flaring simultaneously, focusing on cases where the target source’s flare lasts longer than that of the nearby source. In these cases, the spectral indices remain consistent when comparing the time range in which only the target source is flaring to the full duration of the flare.
@@ 4.1 Spectral hardening in blazar flares
We identified three flares with indications of spectral hardening, all of which fall near or just above our detection threshold (see Table 6). One flare is associated with 4FGL J0238.43116 (1RXS J023832.6311658) with TS, and two with 4FGL J2158.83013 (PKS 2155304), detected independently by the Flip (TS) and Baseline (TS) methods. The flares in PKS 2155304 overlap in time and likely correspond to the same physical event, captured slightly differently due to the distinct selection criteria of the two methods.
@@
Both blazars have been detected by Imaging Atmospheric Cherenkov Telescopes (IACTs): J0238.43116 (Gaté et al. 2017) and PKS 2155304 (Abramowski et al. 2010). The latter has been extensively studied in searches for quasi-periodic oscillations (QPOs; Peñil et al. 2020; Rueda et al. 2022; Peñil et al. 2025; Rico et al. 2025), making it a particularly compelling target for investigating a possible link between spectral hardening and QPOs. Such a connection could help characterize periodic modulations in the jet’s particle acceleration and cooling processes.
@@ 4.2 Physical interpretation
The candidate spectral hardening features identified in this work could, in principle, arise from a limited set of physical processes. In the context of leptonic models, one possibility is a temporary shift of the synchrotron and IC peaks to higher energies, as occasionally suggested during flaring episodes (e.g., Tavecchio et al. 2009; Zeng et al. 2022). Alternatively, a multi-zone scenario may apply, where a secondary SSC component or compact emission region briefly dominates the spectrum (e.g., Abramowski et al. 2012; Abe et al. 2024). A third possibility is that magnetic reconnection or other mechanisms inject freshly accelerated electrons, producing unusually hard particle distributions (e.g., Sahu et al. 2025). While our search among HSP and EHSP blazars finds, at most, very rare and marginal indications of spectral hardening, it is noteworthy that a recent detection of TeV emission from the FSRQ S5 1027+74 revealed an unusually hard -ray spectrum at GeV, illustrating that such extreme features can occasionally arise even outside the HSP population (Paliya et al. 2025).
Hadronic or hybrid scenarios have also been invoked in the literature, where proton synchrotron or photohadronic cascades contribute to -ray emission (e.g., Biteau et al. 2020; Sahu et al. 2025). However, these generally require extreme physical conditions and remain challenging to test with the current data.
Our analysis alone cannot disentangle between these possibilities. The most robust outcome of our study is that such spectral hardening features are extremely rare, with an occurrence rate below 0.1% across thousands of flares. This rarity strongly suggests that the physical conditions required to produce sharp breaks in the GeV band are seldom realized in blazar jets. Establishing which of the above scenarios may apply will require future events detected with higher significance and, crucially, simultaneous multi-wavelength observations.
@@
Our pipeline analyzed 4063 flares and identified two independent flaring episodes with indications of spectral hardening: one in 4FGL J0238.43116 and another in PKS 2155304. The latter is detected with both the Flip and Baseline methods but corresponds to the same flaring activity period. Since both events lie at or just above the detection threshold (TS), the occurrence rate of candidate spectral hardening events is below 0.1%. To our knowledge, there are no theoretical works that provide explicit predictions for the occurrence rate of spectral hardening flares in blazars, as most studies focus on modeling individual events under assumed physical conditions rather than estimating their population frequency (e.g., Tavecchio et al. 2009; Böttcher et al. 2013). This confirms that such features are extremely rare in -ray blazars, if present at all. In the context of standard leptonic models, this rarity suggests that the physical conditions required to produce sharp spectral hardening, such as transitions between distinct electron populations or emission zones, are seldom met. Alternatively, the hardening may emerge only under exceptional jet environments involving magnetic reconnection, multi-zone structures, or hadronic contributions. The scarcity of events reinforces the notion that the dominant blazar emission mechanism is well described by smoothly varying power-law spectra across the Fermi-LAT range, with sharp hardenings representing rare deviations from this behavior.
@@ 5 Summary and conclusions
One of these events is found in 4FGL J0238.43116 and the other two in PKS 2155304, the latter likely corresponding to the same physical flare captured by both selection methods. Given the small number of detections and their moderate significance, these findings are statistically compatible with random fluctuations and do not constitute strong evidence for a widespread occurrence of spectral hardening among blazars. Nonetheless, the identified sources remain of interest for follow-up studies, particularly due to their known HSP frequencies and previous detections at very high energies. Notably, both J0238.43116 and PKS 2155304 have been detected by IACTs. The latter, as a well-studied QPO candidate, stands out as a key target for investigating a possible connection between spectral hardening and QPOs.
Future multi-wavelength observations are essential to further explore these sources, particularly involving IACTs for VHE -ray studies. Our work lays the groundwork for these efforts by identifying key targets and characterizing their tentative spectral hardening behavior. Such observations will be crucial in refining our understanding of the physical processes at play in specific blazar classes and in clarifying the origins of the observed spectral features.
@@ Acknowledgements.
The authors are thankful to D. Yan for private communication on the 1ES 0502+675 analysis. We thank J. Valverde for her helpful revision of the manuscript and J. Ballet, Filippo D’Ammando, and Miguel Ángel Sánchez-Conde for fruitful comments. A. Dinesh and J.L.C. acknowledge the support of MCIN project PID2022-138132NB-C42. A. Domínguez is thankful for the support of the project PID2021-126536OA-I00 funded by MCIN/AEI/10.13039/501100011033. This work was supported by the European Research Council, ERC Starting grant MessMapp under contract no. ; and by the German Science Foundation DFG, research grant ‘Relativistic Jets in Active Galaxies’ (FOR 5195, grant No. 443220636).The Fermi-LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis. These include the National Aeronautics and Space Administration and the Department of Energy in the United States, the Commissariat à l’Energie Atomique and the Centre National de la Recherche Scientifique / Institut National de Physique Nucléaire et de Physique des Particules in France, the Agenzia Spaziale Italiana and the Istituto Nazionale di Fisica Nucleare in Italy, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), High Energy Accelerator Research Organization (KEK) and Japan Aerospace Exploration Agency (JAXA) in Japan, and the K. A. Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden. Additional support for science analysis during the operations phase is gratefully acknowledged from the Istituto Nazionale di Astrofisica in Italy and the Centre National d’Études Spatiales in France. This work was performed in part under DOE Contract DE-AC02-76SF00515.